# The Illusion of Diminishing Returns: Measuring Long Horizon Execution in LLMs

### v1

@@ 5 Discussion
Limitations. As with any “synthetic” task used for a controlled study of LLM capabilities, there are a few limitations of our setup. It does not reflect complexities and sources of error arising in real agentic tasks with a large number of possible actions. In such settings, the number of actions and the accuracy of each action can both vary based on the plan, requiring more careful consideration. It would be interesting future work to study the self-conditioning effect when doing diverse actions instead of repeating the same ones. Our results are observations about pretrained LLMs, and not inherent properties of transformers, so they might change with task-specific finetuning. Improvement on our task is necessary, but not sufficient for long-horizon execution on real-world tasks. Finally, our current task accuracy metric does not account for self-correction. In tasks where mistakes are acceptable and easy to undo, self-correction is a promising direction to improve long-horizon execution.
In the Appendix, we design experiments that dig deeper into our setup. First, in Appendix˜D we note that one step in our task actually requires three operations–retrieval of the value, reading the current state, and adding to it. Individually, we find much better accuracies at long-horizon execution of each of these operations, but taken together, due to increased turn complexity, errors grow much faster. Second, in Appendix˜C we show that the horizon length of different models can vary significantly at different turn complexities. This emphasizes the importance of contextualizing any claims about the length of task a model can complete with the complexity within a turn. Finally, in Appendix˜A we do a preliminary investigation of possible fixes, including self-correction prompting at each turn, and a simple context management technique. Specifically, we find significant improvements from simply removing historic context, which reduces the probability of errors appearing in context for self-conditioning. However, this exploits the Markovian nature of our task and would not work if we added dependencies to arbitrary previous states, such as in dynamic programming (Beniamini et al., 2025).
@@
Outlook. Generative models that maintain accuracy over long horizons will be essential for creating simulated environments (Bruce et al., 2024) to train open-ended agents (Raad et al., 2024). Scaling up the length of tasks a model can complete would be a major step towards realizing the true potential of general agents. By showing that long-horizon execution can be studied on simple tasks, we hope to inspire more research on this capability.
@@ Acknowledgements
We thank Maksym Andriushchenko, Nikhil Chandak, Paras Chopra, Dulhan Jayalath, Abhinav Menon, Sumeet Motwani, Ameya Prabhu, and Shashwat Singh for helpful feedback. AA was funded by the CHIPS Joint Undertaking (JU) under grant agreement No. 101140087 (SMARTY), and by the German Federal Ministry of Education and Research (BMBF) under the sub-project with the funding number 16MEE0444. AA thanks the International Max Planck Research School for Intelligent Systems (IMPRS-IS) and the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. The authors gratefully acknowledge compute time on the Artificial Intelligence Software Academy (AISA) cluster funded by the Ministry of Science, Research and Arts of Baden-Württemberg.
@@
[…]
Appendix A Investigating proposed fixes for long-horizons tasks
A.1 Turn-wise Verification Prompting
We investigate whether self-simulation can be mitigated by explicitly prompting the model to perform active self-correction. At each turn, we instruct the model to first re-validate its previously reported state and, if required, recalculate the full historical sum before processing the current turn’s keys.
The results, shown in Figure 8, are mixed. For the Gemma3 family with CoT, this prompt provides an initial boost in accuracy, successfully breaking the self-simulation loop in early turns. However, the self-verification process significantly increases the number of tokens generated per turn, causing the model to exhaust its context window much sooner, which leads to a sharper performance collapse in later stages. In contrast, the Qwen3 thinking models show negligible improvement. Manually inspecting of their reasoning traces, we find that these models, likely due to their fine-tuning, overthink and frequently fail at the verification step itself, sometimes making arithmetic errors even during their re-calculation process.
@@ A.1 Turn-wise Verification Prompting
Figure 8: Self-verification prompting. Prompting to self-verify does not suffice to fix the self-conditioning effect completely. It leads to overthinking in thinking models and increases the amount of tokens required per turn, leading to faster context consumption in CoT models.
A.2 Context Management
@@
Figure 9: Context Management and Majority Voting on Gemma3 12B. Controlling context size reduces the self-conditioning effect, but relies on the Markovian nature of our task. Majority voting at K=1 provides only minimal improvements over the baseline.
@@ Appendix B Can parallel test-time compute scaling match thinking?
Figure 10: Parallel test time scaling on Gemma3 12B at K=2. Majority voting with the same amount of tokens as CoT traces does not nearly match the performance with CoT.
We also experiment to validate if parallel scaling in test-time compute can achieve the same improvements as thinking. We verify this by testing if parallel majority voting can replicate the gains from either model scale or sequential computation (thinking). To create a fair comparison, we sample multiple outputs from a non-thinking Gemma3 model at each turn, with the number of samples set to match the average token count of its CoT counterpart. The final answer is determined by a majority vote over these parallel generations. From the results in Figure˜10 and 9(b), we see that while majority voting yields a marginal performance improvement, it is insufficient to match the reliability of a larger, non-thinking model, let alone the substantial gains from using CoT reasoning. This suggests that for long-horizon execution, sequential computation provides an advantage that parallel test time scaling cannot match. This contrasts with findings in other domains, such as math or common-sense reasoning, where parallel sampling with self-consistency has been shown to be highly competitive (Snell et al., 2024).
@@ Appendix C Number of turns vs turn complexity
In our experiments, we show that we can increase the length of the task needed to be performed by either (1) increasing the number of turns or (2) increasing the turn complexity, i.e, providing more inputs in the same turn. To investigate the relationship between these two axes, we perform an experiment where a model has to perform a fixed number of operations while varying the turn complexity. A higher turn complexity means the model requires fewer turns to reach the fixed number of operations. Results in Figure 11 indicate there is no strict turn complexity that is consistently the best across model families. Rather, we found that different models behaved quite differently for the same turn complexities. Qwen3 32B seems to show poorer performance at lower turn complexities, indicating that it is unable to perform well over a large number of turns, even if the turns are simple themselves. Gemma3 12B shows a different trend. It reaches accuracy peaks at either extreme of the turn complexity spectrum, failing badly at mid-level turn complexities. This indicates it suffers when the turn complexity and the number of turns are both sufficiently high.
Another axis of evaluating the number of turns vs turn complexity trade-off is the test-time compute used. From an economic view, increasing the number of turns increases the overall cost of inference. We can lower the number of turns by increasing the turn complexity, but that would result in an increase in the per-turn inference cost, as a result of the added complexity. For the same experiment above, we track the number of output tokens used for computation (including thinking tokens) per sample, and again find diverging results for each family.
@@
(b) Average output tokens used to complete the execution vs final accuracy. We see that for Qwen3 32B, more turns lead to more token usage, even at lower turn complexities, pointing to overthinking. Gemma3 12B, on the other hand, uses less tokens for very low turn complexity or very high turn complexity.
@@
Figure 12: Scaling trends hold even enabling Sequential Test Time compute. We compare model performance with thinking disabled (a) against thinking enabled (b, c) at varying turn complexities. (a) Without thinking, all models fail to execute even two steps ($K=2$) in a single turn. (b) In contrast, enabling thinking prevents this performance collapse, with all models successfully handling $K=2$. (c) When the turn complexity is further increased to $K=10$, performance degrades, but a clear scaling trend emerges. (d) This trend is explicitly shown, illustrating that for complex turns, the horizon length increases consistently with model size, reinforcing the benefits of scaling model size even when thinking is enabled.
@@
Appendix D Deconstructing error in retrieve-then-compose
To further isolate the source of execution error, we decompose our task into its two constituent operations–retrieval and addition–and evaluate models on them individually:
  • Retrieval-Only Task. A stateless task where, at each turn, the model is given a key and must simply return the corresponding integer value from the dictionary. No running sum is maintained.
@@ Appendix D Deconstructing error in retrieve-then-compose
  • Prefix-sum Task. A stateful task where, at each turn, the model is given an integer directly and must add it to its previously reported running sum. This isolates the arithmetic and state-tracking component.
@@ E.1 Task Details
We create a dictionary where keys consist of common five-letter English words, and the values consist of integers uniformly sampled from 99-99 to $99$. The range of values is deliberately kept to be large to minimize the chance of an assistant being wrong in an earlier turn correcting its response by pure accident. For our experiment, we first create a fixed set of $100$ keys. Then, we create multiple rollouts (samples) of $50,000$ steps. For each rollout, we uniformly sample a separate set of values to be assigned to each key, in order to increase experimental breadth. Next, at each step, we uniformly sample a key to be provided at that step with replacement. This gives us a list of keys to be processed in order, which is exactly the plan to be executed by the agent.
@@ E.2 Prompting
Each LLM is provided a standardized prompt describing the task at the start of the conversation. This prompt specifies the dictionary containing the five-letter word keys and their corresponding values. Further, the prompt specifies the number of keys that will be provided to the LLM at each subsequent turn. To ensure format following, the prompt also contains few-shot examples of different scenarios. Finally, the LLM is required to provide the running sum after each turn in <answer> tags. An example conversation is shown below.
@@ E.3 Prompting For Thinking Models
To enable models to use chain-of-thought prompting, we add the line ‘‘Think step by step before answering.’’ to the prompt and added CoT traces to the in-context examples. We found that models stop performing CoT reasoning after few turns, as it starts conditioning on the answer format in its history. Thus, we end up including the chain-of-thought trace in the conversation history, to ensure the model does not forget the CoT instruction. This is a trade-off we had to make as it increases the input context of the LLM, however, it was essential to ensure instruction following. Thinking models provided their reasoning in <think> tags, which were removed from the conversation history. No other changes were needed to make the thinking models follow instructions.
@@ E.4 Model Specifications
We use a temperature of $0.6$ and a top-p value of $0.95$ for all Gemma models. For Qwen, we use a temperature of $0.6$ and a top-p value of $0.95$ for thinking mode and a temperature of $0.7$ and a top-p value of $0.8$ for non-thinking as recommended in their documentation. [^1] https://huggingface.co/Qwen/Qwen3-32B
@@ E.5 Compute Details
All experiments were conducted on machines equipped with 4x NVIDIA A100 GPUs with 40/80GB of memory. Frontier model evaluations were performed using OpenRouter.
@@ Appendix F Format Following Failures
In any LLM evaluation, format following failures are a common source of error that is often neglected. In our experiments, any model can have 2 types of format following failures: (1) They do not provide <answer> tags in their answer, and (2) They do not provide a valid integer within <answer> tags. We take multiple steps to minimize format following failures. We ensure clarity in the starting prompt with clear format instructions, as well as few-shot examples. To empirically verify that model errors on our task are actually execution errors and not just format following errors in disguise, for each experiment, we also track the format failure fraction: the fraction of samples that do not correctly follow the format, with the failure being either (1) or (2). It is important to note that while we try to minimize any such error to the best of our abilities, we still count format following as a limitation of the model and hence a source of error.
Our results for format following failures are presented in Figure˜15. For the experiment presented in Figure˜4, we observe that smaller models are more susceptible to format failures, with the Qwen3 family in particular being worse at following format instructions. Overall, the fraction for format following errors is low (around 0.1), with the Qwen3-8B being an exception. We find that the error here actually comes from the model trying to cheat, and do the entire summation inside the <answer> tags (For example, <answer>39 + 51 = 90</answer>). This is explicitly forbidden as we do not allow chain-of-thought or thinking in this experiment, and thus we count this as an error. Gemma3 4B fails at later turns due to a full context, however that does not affect any results, as its accuracies drop much earlier.
Figure 15: Analysis of format following failures. We analyze the fraction of errors attributed to incorrect format following for the experiments presented in Section˜3. Overall, format adherence is high and not the primary source of execution errors.
For the experiments presented in Figure˜15, we find the Qwen3 family to be prone to format following errors in the case where we have thinking disabled for $K=2$. We again find this to be the consequence of models trying to cheat and use extra tokens for computation inside the answer tags. This is fixed by enabling thinking. Following this, the errors in format following become negligible. At $K=10$, we see Gemma3 12B sharply rise to a format failure fraction of 1.0, again due to a full context window.
@@ Appendix G Chain-of-Thought Self-Conditioning
The alternative is to programmatically generate flawed CoT traces. We implemented and experimented with this; however, this introduces its own complexities. For our simple task, there are multiple distinct points of failure within a single trace: an error in the retrieval step (looking up an incorrect value) or an error in the composition step (an arithmetic mistake). A controlled experiment would need to systematically manage the type, frequency, and location of these injected errors, making the setup intractable. Even establishing a “perfectly correct” (Induced Error Rate = 0) baseline history is problematic. A model might have a CoT trace with flawed reasoning (e.g., a minor calculation error that cancels out), which we then replace with the correct final answer. Such a history is also unfaithful.
@@ Appendix H Proof and Analysis of Proposition 1
We can apply this general result to our specific metric, the Effective Task Length ($H_{0.5}$), which is defined as the number of turns at which Task Accuracy drops to $s=0.5$,

### v2

@@ 5 Conclusion
Limitations. As with any “synthetic” task (Allen-Zhu, 2024; Poli et al., 2024; Chollet et al., 2024) used for a controlled study of LLM capabilities, there are a few limitations of our setup. It does not reflect complexities and sources of error arising in real agentic tasks with a large number of possible actions. In such settings, the number of actions and the accuracy of each action can both vary based on the plan, requiring more careful consideration. It would be interesting future work to study the self-conditioning effect when doing diverse actions instead of repeating the same ones. Our results are observations about pretrained LLMs, and not inherent properties of transformers, so they might change with task-specific finetuning. Improvement on our task is necessary, but not sufficient for long-horizon execution on real-world tasks. Finally, our current task accuracy metric does not account for self-correction. In tasks where mistakes are acceptable and easy to undo, self-correction is a promising direction to improve long-horizon execution.
@@
Outlook. Scaling up the length of tasks a model can complete would be a major step towards realizing the true potential of general, open-ended agents (Raad et al., 2024). If they are trained in simulated environments created with generative models (Bruce et al., 2024), maintaining accuracy over a long-horizon becomes doubly important. By showing long-horizon execution can be studied on simple tasks, we hope to inspire more research on this capability, as it is an increasingly important capability in the era of experience (Silver and Sutton, 2025).
@@ Acknowledgements
We thank Maksym Andriushchenko, Nikhil Chandak, Paras Chopra, Dulhan Jayalath, Abhinav Menon, Sumeet Motwani, Mathias Niepert, Lluís Pastor Pérez, Ameya Prabhu, Tim Schneider, Shashwat Singh, and Timon Willi for helpful feedback. AA was funded by the CHIPS Joint Undertaking (JU) under grant agreement No. 101140087 (SMARTY), and by the German Federal Ministry of Education and Research (BMBF) under the sub-project with the funding number 16MEE0444. AA thanks the International Max Planck Research School for Intelligent Systems (IMPRS-IS) and the ELLIS PhD programs for support. The authors gratefully acknowledge compute time on the Artificial Intelligence Software Academy (AISA) cluster funded by the Ministry of Science, Research and Arts of Baden-Württemberg.
@@
Appendix A Investigating Potential Fixes for Self-conditioning
A.1 Turn-wise Self-Verification Prompting
We investigate whether the self-conditioning effect can be mitigated by explicitly prompting the model to perform active self-correction. At each turn, we instruct the model to first re-validate its previously reported state and, if required, recalculate the full historical sum before processing the current turn’s keys.
The results, shown in Figure 8, are mixed. For the Gemma3 family with CoT, this setup provides an initial boost in accuracy, successfully breaking the self-conditioning loop in early turns. However, the self-verification process significantly increases the number of tokens generated per turn, causing the model to exhaust its context window much sooner, which leads to a sharper performance collapse in later stages. In contrast, the Qwen3 thinking models show negligible improvement. Manually inspecting their reasoning traces, we find that these models, likely due to their fine-tuning, overthink and frequently fail at the verification step itself, sometimes making arithmetic errors even during their re-calculation process.
@@ A.1 Turn-wise Self-Verification Prompting
Figure 8: Self-verification does not fix self-conditioning. Prompting to self-verify does not suffice to fix the self-conditioning effect completely. It leads to overthinking in thinking models and increases the amount of tokens required per turn, leading to faster context consumption in CoT models.
A.2 Context Engineering
@@
Figure 9: Context engineering and majority voting on Gemma3 12B. Controlling context size reduces the self-conditioning effect, but relies on the Markovian nature of our task. Majority voting at K=1 provides only minimal improvements over the baseline.
@@ Appendix B Can Parallel Test-time Compute Scaling Match Thinking?
Figure 10: Parallel test time scaling on Gemma3 12B at K=2. Majority voting with the same amount of tokens as CoT traces does not match the performance of CoT.
We also experiment to validate if parallel scaling in test-time compute can achieve the same improvements as thinking. We verify this by testing if parallel majority voting can replicate the gains from either model scale or sequential computation (thinking). To create a fair comparison, we sample multiple outputs from a non-thinking Gemma3 model at each turn, with the number of samples set to match the average token count of its CoT counterpart. The final answer is determined by a majority vote over these parallel generations. From the results in Figure˜10 and 9 (b), we see that while majority voting yields a marginal performance improvement over the base model, it is insufficient to match the reliability of a larger, non-thinking model, let alone the substantial gains from using CoT reasoning. This suggests that for long-horizon execution, sequential computation provides an advantage that parallel test time scaling cannot match. This contrasts with findings in other domains, such as math or common-sense reasoning, where parallel sampling with self-consistency has been shown to be highly competitive (Snell et al., 2024).
@@ Appendix C Number of Turns vs Turn Complexity
In our experiments, we show that we can increase the length of the task needed to be performed by either (1) increasing the number of turns or (2) increasing the turn complexity, i.e, providing more inputs in the same turn. To investigate the relationship between these two axes, we perform an experiment where a model has to perform a fixed number of operations while varying the turn complexity. With a higher turn complexity, the model requires fewer turns to reach the fixed number of operations. Results in Figure˜11 (a) indicate there is no strict turn complexity that is consistently the best across model families. Rather, we found that different models behave quite differently for the same turn complexities. Qwen3 32B seems to show poorer performance at lower turn complexities, indicating that it is unable to perform well over a large number of turns, even if the turns are simple themselves. Gemma3 12B shows a different trend. It reaches accuracy peaks at either extreme of the turn complexity spectrum, failing badly at mid-level turn complexities. This indicates it suffers when the turn complexity and the number of turns are both sufficiently high.
Another axis of evaluating the number of turns vs turn complexity trade-off is the test-time compute used. From a cost perspective, increasing the number of turns increases the overall cost of inference. We can lower the number of turns by increasing the turn complexity, but that would result in an increase in the per-turn inference cost, as a result of the added complexity. For the same experiment above, we track the number of output tokens used for computation (including thinking tokens) per sample, and again find diverging results for each family in Figure˜11 (b).
@@
(b) Average output tokens used to complete the execution vs final accuracy. We see that for Qwen3 32B, more turns lead to more token usage, even at lower turn complexities, pointing to overthinking. Gemma3 12B, on the other hand, uses fewer tokens for very low turn complexity or very high turn complexity.
@@
Figure 12: Scaling trends hold even after enabling Sequential Test Time compute. We compare model performance with thinking disabled (a) against thinking enabled (b, c) at varying turn complexities. (a) Without thinking, all models fail to execute even two steps ($K=2$) in a single turn. (b) In contrast, enabling thinking prevents this performance collapse, with all models successfully handling $K=2$. (c) When the turn complexity is further increased to $K=10$, performance degrades, but a clear scaling trend emerges. (d) This trend is explicitly shown, illustrating that for complex turns, the horizon length increases consistently with model size, reinforcing the benefits of scaling model size even when thinking is enabled.
@@
Appendix D Deconstructing Errors in Retrieve-then-compose Steps
To further isolate the source of execution errors, we decompose our task into its two constituent operations–retrieval and addition–and evaluate models on them individually:
  • Retrieval-Only Task. A stateless task where, at each turn, the model is given a key and must simply return the corresponding integer value from the dictionary. No running sum is maintained. This isolates the retrieval component.
@@ Appendix D Deconstructing Errors in Retrieve-then-compose Steps
  • Prefix-sum Task. A stateful task where, at each turn, the model is given an integer directly and must add it to its previously reported running sum. This isolates the combination of arithmetic and state-tracking components.
@@ E.1 Task Details
We create a dictionary where keys consist of common five-letter English words, and the values consist of integers uniformly sampled from 99-99 to $99$. The range of values is deliberately kept large to minimize the chance of an assistant being wrong in an earlier turn correcting its response by pure accident. For our experiment, we first create a fixed set of $100$ keys. Then, we create multiple rollouts (samples) of $50,000$ steps. For each rollout, we uniformly sample a separate set of values to be assigned to each key, in order to increase experimental breadth. Next, at each step, we uniformly sample a key to be provided at that step with replacement. This gives us a list of keys to be processed in order, which is exactly the plan to be executed by the agent.
@@ E.2 Prompting
Each LLM is provided a standardized prompt describing the task at the start of the conversation. This prompt specifies the dictionary containing the five-letter word keys and their corresponding values. Further, the prompt specifies the number of keys that will be provided to the LLM at each subsequent turn. To ensure format following, the prompt also contains few-shot examples with different turn complexities. Finally, the LLM is asked to provide the running sum after each turn in <answer> tags. An example conversation is shown below.
@@ E.3 Prompting for Thinking Models
To enable models to use chain-of-thought prompting, we add the line ‘‘Think step by step before answering.’’ to the prompt and also add CoT traces to the few-shot examples. We find that models stop performing CoT reasoning after a few turns, as it starts conditioning on the answer format in its history. Thus, we end up including the chain-of-thought trace in the conversation history, to ensure the model does not forget the CoT instruction. This is a trade-off we had to make as it increases the input context of the LLM, however, it was essential to ensure instruction following. Thinking models provided their reasoning in <think> tags, which were removed from the conversation history. No other changes were needed to make the thinking models follow instructions.
@@ E.4 Model Specifications
We use a temperature of $0.6$ and a top-p value of $0.95$ for all Gemma models. For Qwen, we use a temperature of $0.6$ and a top-p value of $0.95$ for thinking mode and a temperature of $0.7$ and a top-p value of $0.8$ for non-thinking as recommended in their documentation. [^1] https://huggingface.co/Qwen/Qwen3-32B We find in Figure˜13 that temperature does not affect the observed trends by much.
@@ E.5 Compute Details
All experiments were conducted on machines equipped with 4x NVIDIA A100 GPUs with 40/80GB memory. Frontier model evaluations were performed using OpenRouter.
@@ Appendix F Format Following Failures
In any LLM evaluation, format following failures are a common source of error that is often neglected. In our experiments, any model can have 2 types of format following failures: (1) They do not provide <answer> tags in their answer, and (2) They do not provide a valid integer within <answer> tags. To minimize format following failures, we ensure clarity in the starting prompt with clear format instructions, as well as few-shot examples. To empirically verify that model errors on our task are actually execution errors and not just format following errors in disguise, for each experiment, we also track the format failure fraction: the fraction of samples that do not correctly follow the format, with the failure being either (1) or (2). It is important to note that while we try to minimize any such error to the best of our abilities, we still count format following as a limitation of the model and hence a source of error.
Our results for format following failures are presented in Figure˜15 for the experiments presented in Section˜3. We observe that smaller models are more susceptible to format failures, with the Qwen3 family in particular being worse at following format instructions. Overall, the fraction for format following errors is low (around 0.1), with the Qwen3-8B being an exception. We find that the error here actually comes from the model trying to cheat, and do the entire summation inside the <answer> tags (For example, <answer>39 + 51 = 90</answer>). This is explicitly forbidden as we do not allow chain-of-thought or thinking in this experiment, and thus we count this as an error. Gemma3 4B fails at later turns due to context length limitations; however, that does not affect any results, as its accuracies drop much earlier.
Figure 15: Format following is not the primary mode of failure. We analyze the fraction of errors attributed to incorrect format following for the experiments presented in Section˜3. Overall, format adherence is high and not the primary source of execution errors.
For the experiments presented in Figure˜15, we find the Qwen3 family to be prone to format following errors in the case where we have thinking disabled for $K=2$. We again find this to be the consequence of models trying to cheat and use extra tokens for computation inside the answer tags. This is fixed by enabling thinking. Following this, the errors in format become negligible. At $K=10$, we see Gemma3 12B sharply rise to a format failure fraction of 1.0, again due to a full context window.
@@ Appendix G Chain-of-Thought Self-Conditioning
Figure 16: CoT does not fix self-conditioning. We observe that even with programmatically generated CoT history, the Gemma3 models cannot mitigate self-conditioning.
The alternative is to programmatically generate flawed CoT traces. We implemented and experimented with this, and as seen in Figure˜16, CoT does not mitigate the self-conditioning effect. However, this setup also introduces its own complexities. For our simple task, there are multiple distinct points of failure within a single trace: an error in the retrieval step (looking up an incorrect value) or an error in the composition step (an arithmetic mistake). A controlled experiment would need to systematically manage the type, frequency, and location of these injected errors, making the setup intractable. Even establishing a “perfectly correct” (Induced Error Rate = 0.00) baseline history is problematic. A model might have a CoT trace with flawed reasoning (e.g., a minor calculation error that cancels out), which we then replace with the correct final answer. Such a history is also unfaithful.
@@ Appendix H Proof and Analysis of Proposition 1
We can apply this general result to our specific metric, the Horizon Length ($H_{0.5}$), which is defined as the number of turns at which Task Accuracy drops to $s=0.5$,